# Schwarzian Theory and Cosmological Constant Problem

### v1

@@ Abstract
as follows from Eq. (9). We define $H_{\Lambda}\equiv\sqrt{\Lambda/3}$ as the characteristic scale of the late-time, $\Lambda$-dominated universe Jazayeri et al. (2016). We will show below that the second term in Eq. (23) is negligible for the present observable universe. Combining Eqs. (23), (24), and (25) then fixes the constant $A=2$. The averaged dark energy density therefore becomes
@@
Since our construction is based on averaging over time reparametrizations, the framework does not privilege any particular time slicing. A time-dependent de Sitter horizon temperature is therefore not suitable. Instead, we introduce a universal, time-independent temperature defined by
$T_{\Lambda}$ $\equiv\frac{H_{\Lambda}}{2\pi}=T_{dS}\,\frac{H_{0}\sqrt{\Omega_{\Lambda}^{0}}}{H(z)}$
$=\frac{T_{dS}}{\sqrt{1+\frac{\Omega_{m}^{0}}{\Omega_{\Lambda}^{0}}(1+z)^{3}+\frac{\Omega_{r}^{0}}{\Omega_{\Lambda}^{0}}(1+z)^{4}+\frac{\Omega_{k}^{0}}{\Omega_{\Lambda}^{0}}(1+z)^{2}}}\,$, (27)
where $H_{\Lambda}$ is defined in Eq. (25). The temperature $T_{\Lambda}$ isolates the $\Lambda$-dependent contribution and coincides with that of a pure de Sitter spacetime. It may be interpreted as the temperature measured by a freely falling observer at the center of the static patch of a pure de Sitter spacetime.
Combining Eqs. (24) and (27) yields
$T_{\Lambda}=\frac{H_{0}\sqrt{\Omega_{\Lambda}^{0}}}{2\pi}\,$, (28)
which is independent of time or redshift. This temperature admits an alternative interpretation. The time-independent $\Lambda$-normalized temperature $T_{\Lambda}$ corresponds to a reference apparent horizon that can be defined at the onset of radiation domination, with radius
$R_{\Lambda}=\frac{1}{H_{0}\sqrt{\Omega_{\Lambda}^{0}}}>\frac{1}{H_{0}}\,.$ (29)
The physical apparent horizon is smaller than this reference scale throughout most of cosmic history and becomes comparable to it only at late times, $z\in[0,\,2]$. In the asymptotic future, the apparent horizon approaches $R_{\Lambda}$, and the spacetime correspondingly approaches pure de Sitter geometry.
@@
From the scaling relation $\rho\propto a^{-3(1+\omega)}$, the equation-of-state parameter for dark energy in this model evolves from the present value $\omega_{\Lambda}=-1$ toward $\omega_{\Lambda}<-1$ at late times. The model therefore predicts a future phantom-like phase that can culminate in a Big Rip singularity, barring additional low-energy phase transitions.

### v2

@@ Abstract
as follows from Eq. (9). Hence, we define $H_{\Lambda}\equiv\sqrt{\Lambda/3}$ as the characteristic scale of the late-time, $\Lambda$-dominated universe Jazayeri et al. (2016). We will show below that the second term in Eq. (23) is negligible for the present observable universe. Combining Eqs. (23), (24), and (25) then fixes the constant $A=2$. The averaged dark energy density therefore becomes
@@
Since our construction is based on averaging over time reparametrizations, the framework does not privilege any particular time slicing. A time-dependent de Sitter horizon temperature is therefore not suitable. Instead, we introduce a universal, time-independent temperature Gibbons and Hawking (1977); Spradlin et al. (2001); Bousso (2002); Cai and Kim (2005); Viaggiu (2016):
$T_{\Lambda}\equiv\frac{H_{\Lambda}}{2\pi}=\frac{H_{0}\sqrt{\Omega_{\Lambda}^{0}}}{2\pi}\,$, (27)
where $H_{\Lambda}$ is defined in Eq. (25). We assume that dark energy and matter are not in thermal equilibrium. The temperature $T_{\Lambda}$ isolates the $\Lambda$-dependent contribution and coincides with that of a pure de Sitter spacetime. It may be interpreted as the temperature measured by a freely falling observer at the center of the static patch of a pure de Sitter spacetime.
Combining Eqs. (24) and (28) yields another equivalent expression:
$T_{\Lambda}$ $=T_{dS}(z)\,\frac{H_{0}\sqrt{\Omega_{\Lambda}^{0}}}{H(z)}$
$=\frac{T_{dS}(z)}{\sqrt{1+\frac{\Omega_{m}^{0}}{\Omega_{\Lambda}^{0}}(1+z)^{3}+\frac{\Omega_{r}^{0}}{\Omega_{\Lambda}^{0}}(1+z)^{4}+\frac{\Omega_{k}^{0}}{\Omega_{\Lambda}^{0}}(1+z)^{2}}}$
$=\frac{T_{dS}(z)}{\sqrt{1+\frac{\Omega_{m}(z)}{\Omega_{\Lambda}(z)}+\frac{\Omega_{r}(z)}{\Omega_{\Lambda}(z)}+\frac{\Omega_{k}(z)}{\Omega_{\Lambda}(z)}}}\,$, (28)
which can be measured at any redshift $z$ but is independent of time or redshift. This temperature admits an alternative interpretation. The time-independent $\Lambda$-normalized temperature $T_{\Lambda}$ corresponds to a reference apparent horizon that can be defined at the onset of radiation domination, with radius
$R_{\Lambda}=\frac{1}{H_{\Lambda}}=\frac{1}{H_{0}\sqrt{\Omega_{\Lambda}^{0}}}>\frac{1}{H_{0}}\,.$ (29)
The physical apparent horizon remains smaller than this reference scale for most of cosmic history and becomes comparable to it only at late times, $z\in[0,\,2]$. In the asymptotic future, the apparent horizon approaches $R_{\Lambda}$, and the spacetime approaches pure de Sitter geometry (see Fig. 1). Thus, $R_{\Lambda}=1/(2\pi T_{\Lambda})$ sets the maximal radius of the physical apparent horizon. Although $T_{\Lambda}$ can be fixed from observational data, a first-principles derivation is not presently available; instead, we interpret it as determined by the maximal size of the observable universe in the asymptotic future. As long as the matter contribution to the cosmic energy budget remains non-negligible relative to dark energy, the apparent horizon radius continues to increase, qualitatively accounting for the smallness of $H_{\Lambda}$, $T_{\Lambda}$, and the observed cosmological constant.
Figure 1: The physical apparent horizon $H(z)^{-1}$ eventually approaches the reference apparent horizon $R_{\Lambda}$ (the red circle).
@@
From the scaling relation $\rho\propto a^{-3(1+\omega)}$, the equation-of-state parameter for dark energy in this model evolves from the present value $\omega_{\Lambda}=-1$ toward $\omega_{\Lambda}<-1$ at late times. The model therefore predicts a future phantom-like phase that can culminate in a Big Rip singularity (see Fig. 2), barring additional low-energy phase transitions. Consequently, $T_{\Lambda}$ does not remain constant at very late times but instead begins to increase, causing $\langle\rho_{\Lambda}\rangle$ to grow more rapidly. In this sense, the initial value given by Eq. (27) represents a lower bound for $T_{\Lambda}$.
Figure 2: In the $\Lambda$-dominated universe, the physical and the reference apparent horizons coincide and eventually both shrink in the far future.